OpenStudy (anonymous):

The cost to produce a custom designed t-shirt is five dollars per shirt plus a 40 dollar design fee. What would be the cost to produce 50 shirts?

OpenStudy (anonymous):

Cost = 5s + 40 --- (s = # of shirts) Cost = 5(50) + 40 Cost = 250 + 40 Cost = $ 290

OpenStudy (anonymous):

Thanks I couldn't figure it out im horrible at math

OpenStudy (anonymous):

no problem :) glad to help